OpenStudy (anonymous):

what is the solution of the system of equations? 2x+2y+3z=-6 3x+5y+4z=3 2x+3y+4z=-10

hartnn (hartnn):

lets try elimination method . subtract 1st equation from 3rd equation, what u get ?

OpenStudy (anonymous):

umm. well i know that the 2x cancel out right?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

now what??

hartnn (hartnn):

no, you had to subtract equations, not just remove 'x'.

hartnn (hartnn):

2x+2y+3z=-6 - 2x+3y+4z=-10 _____________________ 0x - 1y - 1z = 4 got this ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

know what do i do with 2nd equation

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

ok, so y+z=-4 now, 1)multiply 2nd equation with 2, 2)multiply 3rd equation with 3, 3) subtract those two resulting equations, what u get ?

hartnn (hartnn):

yes, so u get 6x +10y +8z = 6

OpenStudy (anonymous):

why am i doing that though

hartnn (hartnn):

so , that you get 6x in both new equations, and when you subtract, 6x gets eliminated! :)

hartnn (hartnn):

yes, correct, now subtract.

hartnn (hartnn):

yes, so now you have 2 equations in terms of y and z y+z=-4 y-4z =36 can you solve these ? its easy ...

hartnn (hartnn):

to find, y and z.

OpenStudy (anonymous):

how do i do that?

hartnn (hartnn):

easy! subtract the equations and y gets eliminated!

OpenStudy (anonymous):

@hartnn now what?

hartnn (hartnn):

z=-8 is correct, now put z=-8 in y+z=-4 and find y.

OpenStudy (anonymous):

is this right

hartnn (hartnn):

yes correct :) now put y=4, z=-8 in any one of the original equations and find x !

OpenStudy (anonymous):

@hartnn is this correct

hartnn (hartnn):

good work! correct :)

OpenStudy (anonymous):

thanks so much!!

hartnn (hartnn):

welcome ^_^